Determine the taxable amount of social security benefits for the following situation.

a. Erwin and Eleanor are married and file a joint tax return. They have adjusted gross income of $41,200, no tax-exempt interest, and $14,420 of Social Security benefits. As a result, $ of the Social Security benefits are taxable.
Answer #1

Solution:

   According to tax laws, For married couple filing a joint tax return, They have to pay taxes on up to 50% of their social security benefits if their combined income is between $32,000 to $44,000. If the combined income is greater than $44,000 they have to pay on up to 85% of their social security benefits.

Combined income = Gross income + 1/2 * Social security benefits.

   = 41200 + 1/2*14420

= 41200 + 7210

   = $48,410.

In this case the combined income is $48,410 which is greater than $44,000. So up to 85% of their social security benefits may be taxable. Total social security benefit is $14420. In that $12000 is subjected to 50% taxable and remaining $2,420 is 85% taxable.

Taxable social security benefits = 50% * 12,000 + 85% *(14420 - 12000)

= 50% * 12,000 + 85% * 2420

= 6,000 + 2,057

= $ 8,057.

Taxable social